1. An ocean wave travels at approximately 1.97 m/s. This is 4 miles per hour. The frequency of the waves is roughly 0.07 hz (or roughly 4 waves per minute). What is the Wavelength?

 

 

  1. Red visible light has a wavelength of 680 nanometers (6.8 x10-7 m). The speed of light is 3.0 x108 m/s. What is the frequency of red visible light?
1
Expert's answer
2021-04-01T18:33:44-0400

(a)

"v=f\lambda,""\lambda=\dfrac{v}{f}=\dfrac{1.97\ \dfrac{m}{s}}{0.07\ Hz}=28.14\ m."

(b)

"c=f\lambda""f=\dfrac{c}{\lambda}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{6.8\cdot10^{-7}\ m}=441.2\ THz."